The demand function for a commodity is Q=6000−30P

Q=6000−30P. Fixed costs are R72000

R72000 and the variable cost is R60 per unit produced. What are the equations for total revenue and total cost in terms of P

P?


1
Expert's answer
2022-02-14T16:11:18-0500

"C=72000+60Q"


Where "C=" Total cost

"Q=" Quantity of units produced


"TR=P\u00d7Q"


Where "TR=" Total revenue

"P=" Price per unit item

"Q=" Number of units sold


but "Q=6000-30P"


"\therefore\>C=72000+60(6000-30P)"

"C=432000-1800P"



"TR=P(6000-30P)"

"=6000P-30P^2"